$SPAD/src/input as-eg2.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)co matops.as
)sh MatrixOpSym
matrix [[3,4],[5,6]]
symmetricPart %

———-

References^{[1] nothing}

4